For the beam and loading shown, and knowing that distance a = 2m, determine the maximum value of the distributed load W so the deflection at midpoint C does not exceed 5 mm Use E = 200 GPa and Ix = 333 x 106 mm4.

D B D А E
Answer #1

Solution :- Deflection at ci using unit load method :- w: KN m D е A B 10 By ) Tourname 2m Dm 12m 12m 10 W=0. - 86 s by=2w. Bunit load at ci- Apply B € ton Dy 2m 2 ന 2m 7 At equili bolum ; By = 0.5; Dy =o. for the Span ED (2: from E): ma=0. for the Sto the indicates that deflection is unit load opposite direction of Given Ac = 5mm 3 5*10 m. E = 200 Gpa = 200*107 W/m2 I= 33